OpenStudy (anonymous):

Solve for x: −2|x − 3| = −12 x = −3, x = 4.5 x = 3, x = 9 x = −3, x = 9 No solutions

OpenStudy (amoodarya):

−2|x − 3| = −12 |x − 3| = −12/-2 |x − 3| = 6 so x-3 =+6 ,-6

OpenStudy (anonymous):

okay thanks homie

OpenStudy (anonymous):

what that makes no sense @amoodarya

OpenStudy (amoodarya):

can you graph that or y=|x-3| ?